a body is vibrating 6000 times in 1 minute if the velocity of sound in air iis 360 m/s find the frecuncy of vibrating and wavelenth of wave produce
 

Grade:9

1 Answers

Arun
5 years ago
Dear Priyansh
 

(a) Frequency of vibration in hertz 

Given, 

Number of vibration in one minute = 6000 

Number of vibrations in one sec = 6000/ 60 =100Hz 

Therefore, Frequency, f = 100 Hz 

(b) Wavelength of the wave produced 

Given, 

Velocity of speed in air, v = 360 ms–1 

Frequency, f = 100 Hz 

Wavelength, λ = ? v = f λ ⇒ λ= 360 ms−1/ 100s−1 = 3.6m 

The wavelength of the wave is 3.6 m

 

Regards

Arun (askIITians forum expert)